OpenStudy (anonymous):

Express the fractions 3/4, 7/16, and 5/8 with the LCD.

OpenStudy (igreen):

LCD is the least common multiple that goes into the denominators of each fraction.

OpenStudy (igreen):

\(\sf \dfrac{3}{\color{red}4}, \dfrac{7}{\color{red}{16}}, \dfrac{5}{\color{red}{8}}\) 4, 16 , 8 Find the multiples: 4: 4, 8, 12, 16, 20, 24 16: 16, 32, 48, 64, 80 8: 8, 16, 24, 32, 40, 48

OpenStudy (igreen):

@hayez

OpenStudy (igreen):

What's the smallest number that all 3 of them have?

OpenStudy (hockeychick23):

3/4, 7/16, and 5/8 3/4*4/4=12/16 7/16 stays the same 5/8*2/2=10/16 Your LCD= 16